OpenStudy (anonymous):

can someone please help me understand this problem. sqareroot 7a +6 squareroot 175a^3 Any and help help would be greatly appreciated

OpenStudy (anonymous):

\[\sqrt{7a}+\sqrt[6]{175a^3}\]

OpenStudy (anonymous):

ok you are not allowed do have common factors for powers in "radicand" and "index"

OpenStudy (anonymous):

here the "index" is 6, the sixth root. and you have an exponent of 3. \[a^3\] you also have an power 2 because \[175=5^2\times 7\] hmm you don't get much out of this, but you do get \[\sqrt[6]{a^3}=\sqrt{a}\] and \[\sqrt[6]{175}=\sqrt[3]{5}\sqrt[6]{7}\]

OpenStudy (anonymous):

oh maybe you are supposed to go backwards? now i am confused. are you supposed to some how add these? because i don't think that you can

OpenStudy (anonymous):

it says to add. and to simplify

OpenStudy (anonymous):

you cannot add because they are not like terms

OpenStudy (anonymous):

you are sure it is 175 rigth? because that is a problem